OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of 3 and contains the point (2, 5).

OpenStudy (bahrom7893):

Okay here's the deal: y-y1 = M(x-x1) plug in 5 for y1 3 for M and 2 x1 and simplify into the form: y = mx+b

OpenStudy (anonymous):

5=6+b?

OpenStudy (anonymous):

no just simplify .. b will be a number!

OpenStudy (bahrom7893):

No: y - 5 = 3(x-2)

OpenStudy (anonymous):

just telll me lyk

OpenStudy (bahrom7893):

Just simplify that, it's not hard.. first multiply out the right side. Then add five to both sides.

OpenStudy (anonymous):

y=3x-9

OpenStudy (bahrom7893):

No.. y-5 = 3(x-2) y-5 = 3x - 6 y = 3x - 6 + 5 y = 3x-1